if a gas exerts a pressure of 102.5 kPa at 25 Kelvin, what is its pressure at a temperature of -100,0C?
1
Expert's answer
2012-06-22T08:40:05-0400
The temperature in Kelvin`s is T2 = 273 - 100 = 173 K
P1/T1 = P2/T2
P2 =
P1 * T2 / T1 = 102.5 * 173 / 25 = 709.3 kPa